# Nonparametric bounds for vaccine effects in randomized trials

### v1

@@ A.1.1 LP-based bounds
Consider the setting of Figures  2 and 3(b)3(d). The following bounds are valid and sharp.
@@ A.1.2 Monotonicity-based bounds
Consider the setting of Figures 2 and 3(b)3(d). Under Assumptions 27(i), and 8(i), the following bounds for the vaccine effect are valid and sharp.
@@ A.2.1 LP-based bounds
Consider the setting of Figure 3(a). The following bounds for the vaccine effect are valid and sharp.
@@ A.2.2 Monotonicity-based bounds
Consider the setting of Figure 3(a). Under Assumptions 5,8(i) and  9(i), the following bounds for the vaccine effects are valid and sharp.
@@ B.1 A proof that Assumption 6 holds under NPSEM-IE model
[…]
@@
[…]
@@
[…]
@@ C.1 Short description of the general LP method
Here, $\mathbf{q}$ denotes the probability vector over the possible response types, $\mathbf{c}^{\top}\mathbf{q}$ encodes the target causal functional, and $\mathbf{R}\mathbf{q}=\mathbf{p}$ captures the linear constraints implied by the observed data. The normalization and nonnegativity constraints ensure that $\mathbf{q}$ defines a valid probability distribution. The optimal value of this linear program yields the sharp lower (or upper) bound for the causal effect of interest under the given assumptions. In the following section, we explicitly construct the vectors $\mathbf{q}$, $\mathbf{c}$, $\mathbf{p}$, and the matrix $\mathbf{R}$ for each setting of interest.
@@ C.3 The target parameter VE, its building blocks and the workflow of the LP method in our context
Our objective is to compute sharp bounds for vaccine efficacy (VE) parameters, such as $VE(0)$. In this section, we detail the derivation of the upper bound for $VE(0)$. The procedures for obtaining the lower bound of $VE(0)$, as well as the bounds for $VE(1)$ and $VE_{T}$, follow analogously.
@@
    We then substitute the extremal values obtained from these LPs into the formula for $VE(0)$. The combinations of these extremal values yield the sharpest (i.e., tightest) possible upper and lower bounds on $VE(0)$ consistent with the data and assumptions.
@@ C.7 Finding the extremal values of VE
C.8 Proof that the LP-based bounds are valid and sharp
The validity of the LP-based bounds follows from the fact that the linear program explores all possible distributions over response types consistent with the observed data and the assumed structural model. By optimizing the target functional over this feasible set, the resulting interval is guaranteed to contain any value compatible with these inputs.
For sharpness, we show that there exists a DGM under which $VE(m)$ for $m\in\{0,1\}$ and $VE_{T}$ attain their respective upper or lower bounds. As in the previous sections, we detail the argument for $VE(0)$, analogous DGMs establish sharpness for $VE(1)$ and $VE_{T}$. We will use the structural equations specified in the NPSEM-IE in Section B.1; here we only modify the equations for $U$, $B^{a,m=-1}$, and $Y^{a,m=0}$ in order to demonstrate sharpness. By Section C.3, we can write
$VE(0)=1-\frac{\pi_{0.1}\,E\!\big(Y^{a=1,m=0}\mid A=1,B^{a=1,m=-1}=0\big)+\pi_{1.1}\,E\!\big(Y^{a=1,m=0}\mid A=1,B^{a=1,m=-1}=1\big)}{\pi_{0.0}\,E\!\big(Y^{a=0,m=0}\mid A=0,B^{a=0,m=-1}=0\big)+\pi_{1.0}\,E\!\big(Y^{a=0,m=0}\mid A=0,B^{a=0,m=-1}=1\big)}\,.$ (19)
Consequently, and using the consistency assumption and that in the trial M=1M=-1 with probability 1, the bounds for $VE(0)$ in Proposition 1 can be written as
$1$- $\frac{\pi_{0.1}\Pr(Y=1\mid A=1,B=0)+\pi_{1.1}}{\pi_{0.0}\Pr(Y=1\mid A=0,B=0)}$ (20)
$\;\leq$ $VE(0)$
$\;\leq$ $1-\frac{\pi_{0.1}\Pr(Y=1\mid A=1,B=0)}{\pi_{0.0}\Pr(Y=1\mid A=0,B=0)+\pi_{1.0}}\,.$
Lower bound is sharp.
Consider the following DGM:
$U$ $\sim\mathrm{Bernoulli}(1/2),$ (21)
$B^{a,m}$ $=m,\;m=\{0,1\}$
$\Pr(B^{a,m=-1}=1)$ $=U,\;a=\{0,1\}$
$\Pr(Y^{a,m}=1|U=u)$ $=aU+\psi_{a}(1-U),\;a,m=\{0,1\}^{2}$
where $0<\psi_{a}<1$ is an arbitrary number. Under randomization and model (21) we have that
$E\!\big(Y^{a=1,m=0}\mid A=1,B^{a=1,m=-1}=1\big)=E\!\big(Y^{a=1,m=0}\mid B^{a=1,m=-1}=1\big)$ $=E\!\big(Y^{a=1,m=0}\mid U=1)=1,$
$E\!\big(Y^{a=0,m=0}\mid A=0,B^{a=0,m=-1}=1\big)=E\!\big(Y^{a=0,m=0}\mid B^{a=0,m=-1}=1\big)$ $=E\!\big(Y^{a=0,m=0}\mid U=1\big)=0$
$E\!\big(Y^{a=1,m=0}\mid A=1,B^{a=1,m=-1}=0\big)=E\!\big(Y^{a=1,m=0}\mid B^{a=0,m=-1}=0\big)$ $=E\!\big(Y^{a=1,m=0}\mid U=0\big)=\psi_{1},$
$E\!\big(Y^{a=0,m=0}\mid A=0,B^{a=0,m=-1}=0\big)=E\!\big(Y^{a=0,m=0}\mid B^{a=1,m=-1}=0\big)$ $=E\!\big(Y^{a=0,m=0}\mid U=0)=\psi_{0},$
In addition, we have that $\psi_{a}=\Pr(Y=1|A=a,B=0)$. Plugging these into the expression for $VE(0)$ in Equation (19) attains the lower bound in Equation (20).
Upper bound is sharp.
We consider the same DGM as in Equation (21) but with different model for $Y^{a,m}$: $\Pr(Y^{a,m}=1|U=u)=(1-a)\mathbb{I}_{\{u=1\}}+\psi_{a}\mathbb{I}_{\{u=0\}},\;m=\{0,1\}.$ Then we get that
$E\!\big(Y^{a=1,m=0}\mid A=1,B^{a=1,m=-1}=1\big)=E\!\big(Y^{a=1,m=0}\mid B^{a=1,m=-1}=1\big)$ $=E\!\big(Y^{a=1,m=0}\mid U=1)=0,$
$E\!\big(Y^{a=0,m=0}\mid A=0,B^{a=0,m=-1}=1\big)=E\!\big(Y^{a=0,m=0}\mid B^{a=0,m=-1}=1\big)$ $=E\!\big(Y^{a=0,m=0}\mid U=1\big)=1$
$E\!\big(Y^{a=1,m=0}\mid A=1,B^{a=1,m=-1}=0\big)=E\!\big(Y^{a=1,m=0}\mid B^{a=0,m=-1}=0\big)$ $=E\!\big(Y^{a=1,m=0}\mid U=0\big)=\psi_{1},$
$E\!\big(Y^{a=0,m=0}\mid A=0,B^{a=0,m=-1}=0\big)=E\!\big(Y^{a=0,m=0}\mid B^{a=1,m=-1}=0\big)$ $=E\!\big(Y^{a=0,m=0}\mid U=0)=\psi_{0},$
Plugging these into the expression for $VE(0)$ in Equation (19) attains the upper bound in Equation (20).
@@ Appendix D Non-linear optimization for bounding VE
bounds_optimized_VE_no_s (available at https://github.com/xlrod1/vaccine_for_publication/tree/main/Non_linear_optimization) to numerically maximize and minimize the VE expressions over the relevant variables. This approach allows optimization of the VE parameters directly and enables the incorporation of additional, easily specified constraints (such as monotonicity or, for example, $\mathrm{E(Y^{a=1,m=0}\mid A=1,B=0)}-\mathrm{E(Y^{a=0,m=0}\mid A=1,B=0)}\leq k$). As mentioned in Section 4.1, this approach is also applicable to computing $VE_{M}(a)$. The optimization was performed using the nloptr and optim packages in R, ensuring robust and flexible computation of sharp bounds under both standard and user-defined constraints.
@@
Appendix E Sharpening the LP-based bounds
In this appendix, we describe how incorporating monotonicity assumptions in either $A$ or $M$ can sharpen the LP-based bounds. First, we present the monotonicity assumptions required to refine the LP-based bounds. These assumptions resemble Assumption 8, but they are formulated at the individual level and hold conditional on $A$ and $B$. Second, we explain how these assumptions are translated into the potential response-type framework and how they are implemented in the LP procedure. Finally, we present the sharpened LP-based bounds corresponding to Propositions 1 and 4, obtained under the monotonicity assumption on $M$. For the accompanying code and illustrative examples demonstrating the resulting bounds under various constraints, see https://github.com/xlrod1/vaccine_for_publication.
E.1 Sharpening the LP-based bounds in Proposition 1
@@ E.2 Implementation in Code
E.3 Sharpened LP-based bounds under Figure 2 and Assumption 10.1
@@
Consider the setting of Figure 2. Under Assumption 10.1, the following bounds for $VE(0)$, $VE(1)$, and $VE_{T}$ are valid and sharp:
@@
E.4 Sharpened LP-based bounds under Figure 3(a)
To sharpen the LP-based bounds in Proposition 4, we extend Assumption 10 to additionally condition on $S$, as follows.
@@ E.4 Sharpened LP-based bounds under Figure 3(a)
Consider the setting of Figure 3(a). Under Assumption 12.1, the following bounds for $VE(0)$, $VE(1)$, and $VE_{T}$ are valid and sharp:
@@ Appendix F A non-LP approach to derive bounds
Now, under the condition in Equation 22, we can replace $E({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Y^{a,m=0}}\mid A=a,B^{a,m=-1}=0,U=u)$ with $E({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Y^{a,m=-1}}\mid A=a,B^{a,m=-1}=0,U=u)$. Thus, we obtain:
@@
Appendix G Proof of Proposition 2: one-sided monotonicity-based bounds under Figure 2
This section is organized as follows. In Section G.1, we present lemmas that will be used to prove the monotonicity-based bounds. In Section G.2, we prove that $E(Y\mid A=a,B=1)\geq E(Y^{a,m=1})$. Based on this result, in Section G.3 we derive the one-sided monotonicity-based bounds. In Section G.4, we prove that these bounds are sharp. Finally, in Section G.5, we derive the monotonicity-based bounds under Assumptions 2 and 7(ii).
G.1 Lemmas
@@
(Esary et al.,, 1967) Let $X=(X_{1},\ldots,X_{n})$ be a multivariate random variable such that each component is independent of the others. Then, for every pair of nondecreasing functions $f$ and $g$,
@@ G.2 Proof that $E(Y\mid A=a,B=1)\geq E(Y^{a,m=1})$
Now, by substituting Equations 25 and 26 into Equation 27, we can show that $E(Y\mid A=a,B=1)\geq E(Y^{a,m=1})$ as follows:
@@
G.3 Proof of the bounds in Proposition 2
@@
G.4 Proof that the bounds in Proposition 2 are sharp
In this section, we prove that the bounds in Proposition 3 are sharp. As demonstrated in Section E, we show that there exists a DGM under which $VE(0)$ attains its lower and upper bounds and $VE_{T}$ attains its lower bound. The DGM is based on Section 2.3, where we review the results from Stensrud et al., (2024), which state that when $U$ is not a common cause of both $B$ and $Y$, then $E(Y^{a,m=1})=E(Y\mid A=a,B=1)$.
For the lower bounds of $VE(0)$ and $VE_{T}$, consider the following DGM, where $U$ does not affect $Y$:
$\Pr(Y^{a,m}=1)$ $=(1-a)\psi_{a}+a,\;m\in\{0,1\},$ (28)
where $\psi_{a}$ is an arbitrary number such that $0<\psi_{a}<1$. Under randomization and model (28), we have
$E(Y^{a=1,m=1})=1,$
$E(Y^{a=1,m=0})=1,$
$E(Y^{a=0,m=0})=\psi_{0}=\Pr(Y=1\mid A=0,B=0).$
In addition, we have $\psi_{a}=\Pr(Y=1\mid A=a,B=0)$. Substituting these into the expressions for $VE(0)$ and $VE_{T}$ yields the lower bounds in Proposition 2.
For the upper bounds of $VE(0)$, consider the following DGM, where $U$ does not affect $Y$:
$\Pr(Y^{a,m}=1)$ $=a\psi_{a}+(1-a),\;m\in\{0,1\},$ (29)
where $\psi_{a}=\Pr(Y=1\mid A=a,B=0)$. Under randomization and model (29), we have
$E(Y^{a=1,m=0})=\psi_{1}=\Pr(Y=1\mid A=1,B=0),$
$E(Y^{a=0,m=0})=1.$
Substituting these into the expression for $VE(0)$ yields the upper bound in Proposition 2.
G.5 The monotonicity-based bounds under Assumptions 2 and 7(ii)
Following the proof in Section G.2, it follows that under Assumption 7(ii), we have:
@@ Appendix H Proof of Proposition 3: two-sided monotonicity-based bounds under Figure 2
Under Assumption 8(i) and the standard assumptions of randomization and consistency, it follows that
@@
H.1 Proof that the bounds in Proposition 3 are sharp
Under the DGM presented in Section G.4, the model for $E(Y^{a,m})$ is not a function of $m$; that is, $E(Y^{a,m})=p_{1.a}$. Therefore, under this DGM, the bounds in Proposition 3 are sharp as well.
H.2 The monotonicity-based bounds under Assumptions 7(ii) and 8(ii)
Under Assumptions 7(ii) and 8(ii), and by incorporating the results from Sections G.5 and H, we get that
@@ Appendix I Proof of Proposition 4: LP-based bounds under Figure 3(a)
I.1 Proof that the LP-based bounds in Proposition 5 are sharp
As in the previous sections, we detail the argument for the lower bound of $VE(0)$; analogous DGMs establish sharpness for the upper bound of $VE(0)$ and for the bounds of $VE(1)$ and $VE_{T}$. To prove the sharpness of the LP-based bounds under Figure 3(a), we consider the following DGM:
$U\sim\mathrm{Bernoulli}(1/2),$ (33)
[…]
$B^{a,m}=m,\;m\in\{0,1\},$
$\Pr(B^{a,m=-1}=1\mid S^{a,m=-1}=s,U=u)=U,$
$\Pr(Y^{a,m}=1\mid U=u)=aU+\psi_{a,s}(1-U),\;a,m\in\{0,1\}^{2},$
where $0<\psi_{a,s}<1$ is an arbitrary number. Under randomization and model (33), we have
$1.\;E$ $\!\big(Y^{a,m=0}|A=a,B^{a,m=-1}=1,S^{a,m=-1}=s\big)$
$=E\!\big(Y^{a,m=0}|B^{a,m=-1}=1,S^{a,m=-1}=s\big)=E\!\big(Y^{a,m=0}|U=1,S^{a,m=-1}=s\big)=a,$
$2.\;E$ $\!\big(Y^{a,m=0}|A=a,B^{a,m=-1}=0,S^{a,m=-1}=s\big)$
$=E\!\big(Y^{a,m=0}|B^{a,m=-1}=0,S^{a,m=-1}=s\big)=E\!\big(Y^{a,m=0}|U=0,S^{a,m=-1}=s\big)=\psi_{a,s}.$
In addition, we have $\psi_{a,s}=\Pr(Y=1\mid A=a,B=0,S=s)$. Substituting these into the expression for $VE(0)$ written using the decomposition in Equation (10) yields the lower bound of $VE(0)$ in Proposition 4.
@@ I.1 Proof that the LP-based bounds in Proposition 5 are sharp
Let $\nu(u)=\frac{\Pr(B=1\mid A=a,U=u,S=s)}{\Pr(B=1\mid A=a,S=s)}$. Then $\nu(u)$ is nondecreasing in $u$ since $\Pr(B=1\mid A=a,S=s,U=u)$ is nondecreasing in $u$, and $E_{U}[\nu(U)]=1$. Thus, by Lemma 1 we have
$\begin{split}E_{U}&\!\left[E\!\left(Y\mid A=a,B=1,U,S=s\right)\nu(U)\right]-E_{U}\!\left[E\!\left(Y\mid A=a,B=1,U,S=s\right)\right]\\ &=\operatorname{Cov}_{U}\!\left(E\!\left(Y\mid A=a,B=1,U,S=s\right),\nu(U)\right)\geq 0.\end{split}$ (36)
Now, by substituting Equations 34 and 35 into Equation 36, we can show that $p_{1.a1s}\geq E\!\left(Y^{a,m=1}\right)$. By using Lemma 2, we can similarly prove the bounds for $E\!\left(Y^{a,m=0}\right)$ and for the discordant-sign case.
@@
If we additionally assume Assumption 8(i), we have
@@
If, alternatively, we assume Assumptions 9(ii) and 8(ii), then
@@
J.3 Proof that the monotonicity-based bounds in Proposition 5 are sharp
The DGM in Appendix H.1 satisfies
$E(Y\mid A=a,B=b,S=s)=E(Y\mid A=a,B=b)\quad\text{for all }a,b,s,$
i.e., $Y\perp\!\!\!\perp S\mid(A,B)$. Therefore, the bounds in Proposition 5 coincide with those in Proposition 2. Since Appendix E constructs DGMs that attain both endpoints of the VEs, these bounds are attained under the present DGM and are therefore sharp.
@@
    We then substitute the extremal values from these LPs into the formula for $VE(0)$. The combination of these extremal values yields the sharpest (i.e., tightest) possible upper and lower bounds on $VE(0)$ consistent with the data and assumptions.
@@
K.1 Proof that the LP-based bounds in Proposition 6 are sharp
Since:
  1. (i)
    Figure 2 is a special case of Figures 3(b)3(d), obtained by setting to zero some of the arrows $S\rightarrow Y$, $B\rightarrow Y$, and $U\rightarrow S$ in Figures 3(b)3(d).
  2. (ii)
    The analytical bounds under these figures coincide with those derived for Figure 2.
It follows that, because the bounds under Figure 2 are sharp, the bounds under Figures 3(b)3(d) are sharp as well.

### v2

@@ A.1.1 LP-based bounds
Consider the setting of Figures  2 and 3(b)3(d). The following bounds hold.
@@ A.1.2 Monotonicity-based bounds
Consider the setting of Figures 2 and 3(b)3(d). Under Assumptions 271, and 81, the following bounds for the vaccine effect hold.
@@ A.2.1 LP-based bounds
Consider the setting of Figure 3(a). The following bounds for the vaccine effect hold.
@@ A.2.2 Monotonicity-based bounds
Consider the setting of Figure 3(a). Under Assumptions 5,81 and  91, the following bounds for the vaccine effects hold.
@@ B.1 A proof that Assumption 6 holds under NPSEM-IE model
B.2 A proof that $Y_{i}^{a,m}\perp_{\mathcal{T}_{VI}}A_{i},S_{i}^{a,m=-1}$ under NPSEM-IE model and Figure 3(a)
@@
where Step 1 follows from the fact that under Assumption 1 for $m\in\{0,1\}$, $B_{i}^{a^{\prime},m}=m$. Step 2 follows since $A_{i}\perp U_{i},\varepsilon_{Y_{i}}|S_{i}^{a^{\prime},m=-1}$. Step 3 follows since $S_{i}^{a^{\prime},m=-1}=f_{S}(a^{\prime},\varepsilon_{S_{i}})$, $U_{i}=f_{U}(\varepsilon_{U_{i}})$ and thus $\varepsilon_{S_{i}}\perp(\varepsilon_{Y_{i}},\varepsilon_{U_{i}})$.
@@
where Step 1 follows from the fact that (I) under Assumption 1, for $m\in\{0,1\}$, $B_{i}^{a^{\prime},m}=m$ and since we conditioned on $B_{i}^{a^{\prime},m=-1}=m$, and (II) $\Pr(S_{i}^{a^{\prime},m=-1}=S_{i}^{a^{\prime},m})=1$, and since we conditioned on $S_{i}^{a^{\prime},m=-1}=s$. Step 2 follows since $S_{i}^{a^{\prime},m=-1}|\{U_{i}=u\}=h_{S}(a^{\prime},u,\varepsilon_{S_{i}})$, $A_{i}=f_{A}(\varepsilon_{A_{i}})$ and thus $A_{i},S_{i}^{a^{\prime},m=-1}\perp\varepsilon_{Y_{i}}|B_{i}^{a^{\prime},m=-1},U_{i}=u$. Step 3 follows since $B_{i}^{a^{\prime},m=-1}\perp\varepsilon_{Y_{i}}\mid U_{i}=u$. Step 4 follows since we assume that the only random component, $\varepsilon_{Y_{i}}$, is the same for each individual $i$.
@@ C.1 Short description of the general LP method
Here, $\mathbf{q}$ denotes the probability vector over the possible response types, $\mathbf{c}^{\top}\mathbf{q}$ encodes the target causal functional, and $\mathbf{R}\mathbf{q}=\mathbf{p}$ captures the linear constraints implied by the observed data. The normalization and nonnegativity constraints ensure that $\mathbf{q}$ defines a valid probability distribution. The optimal value of this linear program yields the lower (or upper) bound for the causal effect of interest under the given assumptions. In the following section, we explicitly construct the vectors $\mathbf{q}$, $\mathbf{c}$, $\mathbf{p}$, and the matrix $\mathbf{R}$ for each setting of interest.
@@ C.3 The target parameter VE, its building blocks and the workflow of the LP method in our context
Our objective is to compute bounds for vaccine efficacy (VE) parameters, such as $VE(0)$. In this section, we detail the derivation of the upper bound for $VE(0)$. The procedures for obtaining the lower bound of $VE(0)$, as well as the bounds for $VE(1)$ and $VE_{T}$, follow analogously.
@@
    We then substitute the extremal values obtained from these LPs into the formula for $VE(0)$. The combinations of these extremal values yield the upper and lower bounds on $VE(0)$ consistent with the data and assumptions.
@@ Appendix D Non-linear optimization for bounding VE
bounds_optimized_VE_no_s (available at https://github.com/xlrod1/vaccine_for_publication/tree/main/Non_linear_optimization) to numerically maximize and minimize the VE expressions over the relevant variables. This approach allows optimization of the VE parameters directly and enables the incorporation of additional, easily specified constraints (such as monotonicity or, for example, $\mathrm{E(Y^{a=1,m=0}\mid A=1,B=0)}-\mathrm{E(Y^{a=0,m=0}\mid A=1,B=0)}\leq k$). As mentioned in Section 4.1, this approach is also applicable to computing $VE_{M}(a)$. The optimization was performed using the nloptr and optim packages in R, ensuring robust and flexible computation of the bounds under both standard and user-defined constraints.
@@
Appendix E Obtaining narrower LP-based bounds
In this appendix, we describe how incorporating monotonicity assumptions in either $A$ or $M$ can be used to obtain narrower LP-based bounds. First, we present the monotonicity assumptions required to refine the LP-based bounds. These assumptions resemble Assumption 8, but they are formulated at the individual level and hold conditional on $A$ and $B$. Second, we explain how these assumptions are translated into the potential response-type framework and how they are implemented in the LP procedure. Finally, we present the resulted LP-based bounds corresponding to Propositions 1 and 3, obtained under the monotonicity assumption on $M$. For the accompanying code and illustrative examples demonstrating the resulting bounds under various constraints, see https://github.com/xlrod1/vaccine_for_publication.
E.1 Refining the LP-based bounds in Proposition 1
@@ E.2 Implementation in Code
E.3 Refined LP-based bounds under Figure 2 and Assumption 11.1
@@
Consider the setting of Figure 2. Under Assumption 11.1, the following bounds for $VE(0)$, $VE(1)$, and $VE_{T}$ hold.
@@
E.4 Refined LP-based bounds under Figure 3(a)
o obtain LP-based bounds that are narrower than those in Proposition 1, we extend Assumption 11 to additionally condition on $S$, as follows.
@@ E.4 Refined LP-based bounds under Figure 3(a)
Consider the setting of Figure 3(a). Under Assumption 13.1, the following bounds for $VE(0)$, $VE(1)$, and $VE_{T}$ hold.
@@ Appendix F A non-LP approach to derive bounds
Now, under the condition in Equation (19), we can replace $E({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Y^{a,m=0}}\mid A=a,B^{a,m=-1}=0,U=u)$ with $E({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Y^{a,m=-1}}\mid A=a,B^{a,m=-1}=0,U=u)$. Thus, we obtain:
@@
Appendix G Proof of the one-sided monotonicity-based bounds under Figure 2
This section is organized as follows. In Section G.1 we provide the one-sided one-sided monotonicity-based bounds under Figure 2. In Section G.2, we present lemmas that will be used to prove the monotonicity-based bounds. In Section G.3, we prove that $E(Y\mid A=a,B=1)\geq E(Y^{a,m=1})$. Based on this result, in Section G.4 we derive the one-sided monotonicity-based bounds. Finally, in Section G.5, we derive the monotonicity-based bounds under Assumptions 2 and 72.
G.1 Proposition: one-sided one-sided monotonicity-based bounds under Figure 2
Proposition 13.
Consider the setting of Figure 2. Under Assumptions 2 and 71, the following bounds for $VE(0)$ and $VE_{T}$ hold.
$1-\frac{1}{p_{1.00}}$ $\leq VE(0)\leq 1-p_{1.10},$
$1-\frac{p_{1.11}}{p_{1.00}}$ $\leq VE_{T}\leq 1.$
Under Assumptions 2 and 7 2, the bounds for $VE(0)$ and $VE_{T}$ given in Appendix G hold.
G.2 Lemmas
@@
(Esary et al., 1967) Let $X=(X_{1},\ldots,X_{n})$ be a multivariate random variable such that each component is independent of the others. Then, for every pair of nondecreasing functions $f$ and $g$,
@@ G.3 Proof that $E(Y\mid A=a,B=1)\geq E(Y^{a,m=1})$
Now, by substituting Equations 22 and 23 into Equation (24), we can show that $E(Y\mid A=a,B=1)\geq E(Y^{a,m=1})$ as follows:
@@
G.4 Proof of the bounds in Proposition 13
@@
G.5 The monotonicity-based bounds under Assumptions 2 and 72
Following the proof in Section G.3, it follows that under Assumption 72, we have:
@@ Appendix H Proof of Proposition 2: two-sided monotonicity-based bounds under Figure 2
Under Assumption 81 and the standard assumptions of randomization and consistency, it follows that
@@
H.1 The monotonicity-based bounds under Assumptions 72 and 82
Under Assumptions 72 and 82, and by incorporating the results from Sections G.5 and H, we get that
@@ Appendix I Proof of Proposition 3: LP-based bounds under Figure 3(a)
Let $\nu(u)=\frac{\Pr(B=1\mid A=a,U=u,S=s)}{\Pr(B=1\mid A=a,S=s)}$. Then $\nu(u)$ is nondecreasing in $u$ since $\Pr(B=1\mid A=a,S=s,U=u)$ is nondecreasing in $u$, and since $U\perp A,S$, we have
$E_{U}[\nu(U)]=E_{U|A,S}[\nu(U)]=1.$
Thus, by Lemma 1 we have
$\begin{split}E&\!\left(Y\mid A=a,B=1,S=s\right)-E\!\left(Y^{a,m=1}\right)\\ =&E_{U}\!\left[E\!\left(Y\mid A=a,B=1,U,S=s\right)\nu(U)\right]-E_{U}\!\left[E\!\left(Y\mid A=a,B=1,U,S=s\right)\right]\\ =&E_{U}\!\left[E\!\left(Y\mid A=a,B=1,U,S=s\right)\nu(U)\right]-E_{U}\!\left[E\!\left(Y\mid A=a,B=1,U,S=s\right)\right]E_{U}[\nu(U)]\\ =&\operatorname{Cov}_{U}\!\left(E\!\left(Y\mid A=a,B=1,U,S=s\right),\nu(U)\right)\geq 0.\end{split}$ (30)
Now, by substituting Equations 28 and 29 into Equation (30), we can show that $p_{1.a1s}\geq E\!\left(Y^{a,m=1}\right)$. By using Lemma 2, we can similarly prove the bounds for $E\!\left(Y^{a,m=0}\right)$ and for the discordant-sign case.
@@
If we additionally assume Assumption 81, we have
@@
If, alternatively, we assume Assumptions 92 and 82, then
@@
    We then substitute the extremal values from these LPs into the formula for $VE(0)$. The combination of these extremal values yields the upper and lower bounds on $VE(0)$ consistent with the data and assumptions.